OpenStudy (anonymous):

Find the 9th term of the geometric sequence with a2 = 0.3 and a6 = 0.00003.

OpenStudy (nikita2):

an = a0*q^n. q^4 = a6/a2 = 0.00001 -> q = 0.1. And a9 = a6*q^3 = 0,000000003